FDC Weekly Meeting
Date: January 7, 2010
Participants: Daniel, Simon, Beni, Eugene, Lubomir, Mark S., Bill, Roger, Casey, Kim, Fernando, Glenn, Chuck
Next Meeting: January 14, 2010
- The new (new) FDC group leader will be Lubomir Pentchev (Hall D staff scientist). Lubomir just started in Hall D this week and is now working to get up to speed on all things Hall D and all things FDC as we move into construction. - Beni will be focussing his efforts on the CDC, however, he will still have significant short-term participation in FDC work and will help to lead the work on the test plan for the full-scale FDC prototype. - Eugene is working to hire an FDC work coordinator to lead the construction efforts. The posting is working its way through lab management. Eugene expects that this hire should be made before the end of March.
General Construction Updates
- Pre-break status: With the full-scale prototype in the clean tent mounted on the assembly fixture, we were able to flow gas through all three layers at a rate of 150 cc/min. The leak rate in each individual layer was about the same. The detector was installed into the cosmic ray facility. Gas was flowed and leak checking began. We established gas flow at 200 cc/min and a few new leaks were detected near the gas inlet/outlet of each layer and at the two points where the rail mounts were attached. - Present status: We started up again after the break and found that we could not get one of the detector layers to flow. In addition, we confirmed that there were leaks opening and closing as we rotated the detector from vertical to horizontal. Bottom line, the half package was too floppy and the small o-rings just were insufficient to make up for the flexing. - We decided to open up the detector and replace the o-rings with a larger size stock. The softness of the foam core/rubber-coated material is essentially the same as what we had before, but the size is now 0.093 in instead of 0.070 in. In our work on the spacers, we also found that the Kapton tubes used on the gas inlets/outlets had become delaminated and were leaking badly. We redesigned them, replacing the Kapton with a short Nylon piece. We modified how these gas tubes were attached to the spacer ring to reduce flow impedance. - So far we have completed a single layer test and were satisfied that the new o-ring and inlet/outlet configuration (and not using any axle grease) is at least as good as what we had before (pre-break) on the assembly fixture. We may add some new stiffening members to the external compression rings to help reduce the flexing of the full-scale prototype when installed in the cosmic ray facility. - On inspection of the single layer assembly, we believe that we have found general leakiness in the regions where there are vias. We sealed up the potential problem regions on the first layer with Humaseal. We will also inspect and seal potentially problematic regions on the second and third wire boards as well. - We should have the detector back together this week and installed into the cosmic ray test stand. We will then go through a round of leak studies. Hopefully early next week the detector will be ready to start on the test plan. - Mark has been tried to find some existing equipment to measure oxygen contamination in our exhaust gas. Nothing has been found. He will keep trying. We may ultimately have to send samples out for testing. - The preliminary FDC stack assembly document draft is located at: /u/group/halld/Engineering/PRELIMINARY DOCUMENTS.
Frame Redesign Studies
- Simon has completed work on a new geometry file for the Monte Carlo. Beni has contacted Mihajlo Kornicer at IU to have him study the impact of this geometry on photon conversions similar to what they did for the chamber review two years ago. We should at least know shortly how long this study will take to complete. - Bill has completed work on the drawings for the new frame G10 support pieces for the wire and cathode frames. These pieces are actually the same for both frame types. The drawings were submitted to the machine shop before the break and the parts should be ready in about a week or so. We plan to modify the 5th wire frame using the solder sample boards with the new design. For the cathodes, Bill would like to put together a cathode sandwich with dummy cathodes. We will consider if there are other suitable lighter weight materials that G10 as we move forward. Stay tuned for details.
Wire Frame Update
- The wire frame construction document is located at: /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. - The parts of the jig to use for the wire electroplating operation are in our hands. They will be assembled as we find time. - We need to prepare a document for the electrical QA of the preamp connectors on the wire boards. - Kim is following up with the contacts regarding large-area PCBs. She has placed calls, but has not heard back from the board supplier or the board house. Kim will follow up with these companies to let them know of our urgency in pursuing this option. Stay tuned. - DSC and Eugene met with Kim's supervisor about an agreement on Kim's time for the wire frame redesign. She will be starting on this work full time in 3 weeks and then basically devote 100% of her work time to completing the redesign. It looks like we should be able to begin wire frame PCB procurement in March with frame construction starting in roughly the same time scale. DSC, Kim, and Fernando will have some meetings over the next three weeks to go over the design work in more detail so that Kim can get started quickly when she can shift over to the board redesign. - Plans for Phase 1 and Phase 2 soldering: The Phase 1 soldering will be done at an outside company. Fernando thinks this is straightforward work if our boards are the size they currently are. If we have single piece STBs and HTVBs, then we will have to do some digging. The Phase 2 work will have to be done onsight. We will have to think about lining up the skilled manpower for this work as we move into the new year.
- The current draft of the cathode construction document is located at: /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. - We need to prepare a document for QA/stuffing/cleaning for the cathode boards and a similar document for the cathode daughter boards and ground boards. Casey has a draft of a document that Mark S. began preparing. He will take this over and complete the document. - The cathode material order should be in shipment to the U.S. from Japan now. It is being sent to the U.S. distributor in Mn. They will inspect the order. If Allflex (only some 30 miles away) will agree to store the material until the start of the board work, we thought that this would be best. - Before the break DSC, Bill, and Roger met with regard to final design of the cathode boards and rigid-flex assemblies. We discussed the plans from that meeting. We had thought about modifying the ground layout to increase the trace-to-trace pitch, but Fernando said that this was a bad idea from a signal-to-noise standpoint. Instead, he embraced the idea of increasing the trace-to-trace pitch between signal lines and the isolating ground lines both on the cathode board itself and the rigid-flex assembly. - Fernando suggested that the half-strips in the middle of the board be made with square ends as opposed to rounded ends. DSC will add this to Roger's work list. - We would like to push to get the cathode board procurement done on the time scale of this March. Stay tuned. - DSC talked to David Lawrence about finalizing the Monte Carlo studies to dictate the size of the cathode beam hole (where the copper strips are remove). He thinks that this work should be completed on the time scale of 4 to 6 weeks. One possible hold up is consideration of running the solenoid at reduced field. Stay tuned.
Cosmic Ray Test Stand
- The current version of the full-scale prototype test plan is on the GlueX portal as (GlueX-doc-1304). - A while ago we focussed on the wiggles seen in the reconstructed wire position from the cathodes (see log entry 419) in the FDC logbook. We believe that the periodicity seen is due to the algorithm Beni is using to determine the U and V centroids. Stay tuned. - To connect the full-scale prototype to the readout electronics, Fernando will design a simple translator board that allows connection from the 24-channel preamp connector to the 16-channel connectors on the ADCs and TDCs. These boards should be ready in early February. - Beni believes that we have all of the HV and LV cables needed for the full-scale prototype tests. We have a few signal cables to make and we are rounding up the parts and pieces now. - To enable the full-scale prototype detector to fit better into the cosmic ray test stand, we had to do some minor repositioning/reconfiguring of the IUCF aluminum chambers. Beni is now reconnecting the chambers. - Mark S. will work with Beni to put together an adjustment system to position the source and trigger scintillator with simple horizontal and vertical adjustements. - Beni stated that the readout tests of the full-scale prototype should be able to commence next week.
- Update on wire sample measurements - A work plan has been set up and the work is set to begin shortly. DSC will work to get an update for our next meeting. - We should have the full amount of Rohacell in hand for production frame construction. We will need to place a G10 order. Bill estimates a 2 to 3 month lead time on this order. We obviously need to get on this as soon as possible (once we figure out what frames we are going to construct). - Matt Shepherd will be meeting with IUCF Director Jim Musser on Friday to discuss the status of our approach to Phase 3 wire winding. Stay tuned. - Next IUCF meeting on Phase 3 preparations will take place in about 3 weeks. See the work plan at: Phase 3 Preparation Work. - Master list on wire and cathode frame redesign work is located at: Frame Modification Work.
- The FDC short-term work list has been posted on the FDC web site. This is continually being updated and DSC welcomes any feedback or comments from the group.
Minutes prepared by Daniel. Send any comments or corrections along.